OpenStudy (anonymous):

HELP ME AND WE CAN GET MARRIED (btw will totally except direct answers) :3 What is 2.48 expressed as the quotient of two integers in simplest form? is it 12/25

OpenStudy (anonymous):

2 12/25

OpenStudy (anonymous):

what you got is correct but in the future, to correct your answer you multiply the whole number by the number in the denominator so (2x25) and add the number in the number in the numerator which will be (2x25+12) all over the number in the denominator so you get (2x25) = 50+12 = 62/25 and you divide the number and get 2.48

OpenStudy (perl):

$$ \Large 2.48 = \frac {2.48}{1} = \frac {2.48}{1} \times \color{red}{\frac{100}{100}} = \frac{248}{100} $$

OpenStudy (perl):

you can reduce that to lowest terms

OpenStudy (anonymous):